OpenStudy (anonymous):

How would you use the quadratic formula to solve x^2+5x=-2. Using compete sentences.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

maybe we can work the prob using math, and you do the sentences part alone?

OpenStudy (anonymous):

Okay

OpenStudy (solomonzelman):

when the equation is \(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\). you use the following formula: \(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightcyan ,border:2px solid black ]{ \LARGE{x=~} \huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}}{2 \color{blue}{a}} }~ }}}\) (to transfer your equation to the needed \(\large \color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\) form, you have to add 2 to both sides.)

OpenStudy (anonymous):

So it will be x^2+5+2=0

OpenStudy (solomonzelman):

what do you get after adding 2 to both sides?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\) ` ... comparing ... ` \(\LARGE\color{black}{ \color{blue}{1} x^2+ \color{magenta}{5}x+ \color{red}{2}=0 }\)

OpenStudy (solomonzelman):

Now, please plug in the a b c for me....

OpenStudy (anonymous):

What do you mean by plug in?

OpenStudy (anonymous):

I have no clue how to do this. I am just learning it

OpenStudy (usukidoll):

she meant let a = 1, b = 5, and c = 2 we will have this equation \[x^2+5x+2=0\]

OpenStudy (anonymous):

Okay and then what do I do from there?

OpenStudy (usukidoll):

Well, to solve this faster we need the discriminant \[b^2-4ac\] so let b = 5, a =1, and c =2

OpenStudy (solomonzelman):

who is she?

OpenStudy (usukidoll):

umm I thought that was you?

OpenStudy (solomonzelman):

well, if you are thinking of me as of a good looking girl, rather than a bad looking girl, then you in a sense did tackle my identity.

OpenStudy (usukidoll):

anyway plug in b =5, c =2, and a = 1 into the discriminant formula \[b^2-4ac\] @J_slate23

OpenStudy (usukidoll):

as in replace the b with 5, replace the c with 2 and replace the a with 1

OpenStudy (anonymous):

So 5^2=4(1)(2)?

OpenStudy (usukidoll):

yes but replace the = with -

OpenStudy (usukidoll):

\[5^2-4(1)(2)\]

OpenStudy (anonymous):

So 5^2-4(1)(2)?

OpenStudy (usukidoll):

mhm so now what's 5^2 and what's 4(1)(2)

OpenStudy (solomonzelman):

\(\normalsize\color{ slate }{\Huge{\bbox[5pt, lightyellow ,border:2px solid black ]{ \LARGE{x=~} \huge{ \frac{-\color{magenta}{5} \pm\sqrt{ \color{magenta}{5} ^2-4 \color{blue}{(1)} \color{red}{(2)}}}{2 \color{blue}{(1)}} }~ }}}\) and your equation is \(\LARGE\color{black}{ \color{blue}{1} x^2+ \color{magenta}{5}x+ \color{red}{2}=0 }\).

OpenStudy (anonymous):

25 and 8

OpenStudy (usukidoll):

yes so now what's 25-8

OpenStudy (anonymous):

17

OpenStudy (usukidoll):

yes! so for the bottom denominator what is 2(1) ?

OpenStudy (anonymous):

2

OpenStudy (usukidoll):

yes

OpenStudy (usukidoll):

so now you have \[\frac{5 +\sqrt{17}}{2}, \frac{5-\sqrt{17}}{2}\] I split the signs up because the latex for the quadratic formula is big and nasty

OpenStudy (usukidoll):

OH MAN I forgot the -!

OpenStudy (usukidoll):

\[\frac{-5 +\sqrt{17}}{2}, \frac{-5-\sqrt{17}}{2}\]

OpenStudy (anonymous):

Is that the answer? If so that was really easy

OpenStudy (usukidoll):

yeah.. I just split the sign in the middle up. the latex for the entire quadratic equation with + and - together was a pain

OpenStudy (usukidoll):

and 17 is not a perfect square... so we can't simplify .. not to mention 17 is a prime

OpenStudy (anonymous):

So the final answer is -5+-sqrt17/2?

OpenStudy (usukidoll):

yeah

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

That is easier then I expected

OpenStudy (anonymous):

Can you check my math for another one? @UskiDoll

OpenStudy (usukidoll):

sure

OpenStudy (anonymous):

Solve x^2+4x-12=0 by completing the square. X^2+4x-12=0 Add 12 X^2+4x=12 (X+2)^2-4=12 (x+2)^2=16 (x+2)= 4 or -4 X= 6 or -6 is my answer

OpenStudy (anonymous):

Is that correct @UskiDoll

OpenStudy (usukidoll):

\[x^2+4x-12=0 \] \[x^2+4x=12 \] \[4 \times \frac{1}{2} = \frac{4}{2} = 2\] \[2^2 = 4\] \[x^2+4x+4=12+4 \] \[x^2+4x+4=16 \] \[(x+2)^2 = 16 \] square root on both sides \[\sqrt{(x+2)^2} = \sqrt{16}\] \[x+2=4, x+2=-4\] \[x=2, x=-6\]

OpenStudy (usukidoll):

x+2 = 4 on your part is wrong. subtract 2 from both sides, but -6 is correct

OpenStudy (anonymous):

Okay. Thank you

OpenStudy (anonymous):

plug it in -b+_ b2-4ac \[\sqrt{b2-4ac}\] / 4a good luck

OpenStudy (usukidoll):

The problem is already done ^_^